# Multimodal ArXiv: A Dataset for Improving Scientific Comprehension of Large Vision-Language Models

### v2

@@ 4.1.3 Analysis
We investigate how different subject domains affect mathematical reasoning ability using pairs of questions and answers (QA). We focus on six domains with more than 5K samples each. From each domain, we randomly choose a subset of 5K samples to ensure fairness in comparison. We then fine-tune the Qwen-VL-Chat base model using QA pairs from each domain and observe how it affects the model’s accuracy compared to its original state. Figure 4 demonstrates the relative accuracy changes (i.e., Accuracy after Fine-tuningOriginal $\frac{\text{Accuracy after Fine-tuning}}{\text{Original Accuracy}}-1$) after training the model on QA pairs from each domain. Our findings reveal several key points: (i) QA pairs from the Computer Science (CS) domain are highly effective for improving mathematical reasoning ability, achieving a notable 27.09% relative improvement. We attribute this to the compositional nature the the CS area. (ii) The most beneficial domain varies depending on the specific task. For instance, QA pairs from astrophysics domains enhance geometry problem-solving, while those from Condensed Matter improve performance in math word problems. (iii) Most domains hurt the Figure QA task. This suggests that synthetic Figure QA might not be the best benchmark for assessing realistic reasoning ability. These findings underscore the efficacy of generated QA pairs and offer valuable insights for future research, such as adjusting task-specific weights in the dataset accordingly.
@@ 4.2.1 Evaluated Tasks
We design four vision-to-text tasks to benchmark LVLMs’ ability to comprehend scientific figures. Single-Figure Captioning Similar to the traditional image captioning setup (Lin et al., 2014; Hsu et al., 2021), single-figure captioning requires the model to generate a caption for the given single figure. The captions generated by the model are expected to encapsulate the nuanced details within these figures, including numbers and mathematical formulas, presenting a unique challenge for models to identify and articulate these elements accurately. Formally, given an image-caption pair $(I,C)$, the LVLM $\mathcal{M}$ is asked to generate the caption given an instruction prompt $P_{s}$ to hint the goal of scientific captioning:
$\hat{C}=\mathcal{M}(I,P_{s}),$ (1)
@@
Multiple-Figure Captioning We introduce a more intricate challenge involving applying reasoning across multiple images. This task, termed Multiple-Figure Captioning, necessitates the model to craft a comprehensive summary caption for subfigures. As exemplified in Figure 2, the model is tasked with generating an overarching caption for two or more subfigures, leveraging visual clues to draw comparisons and formulate summary captions. Formally, given a list of figures $L=\left(I_{1},\ldots,I_{n}\right)$, the model is to generate the ground-truth main caption $C$ by considering all the semantics in the figures with a task prompt $P_{m}$:
$\hat{C}=\mathcal{M}(L,P_{m})=\mathcal{M}(I_{1},\ldots,I_{n},P_{m}).$ (2)
Contextualized Image Captioning Inspired by the evolving in-context learning capabilities of LLMs (Brown et al., 2020b; Dong et al., 2022), we introduce a contextualized captioning task to examine the in-context learning ability of LVLMs. In this task, the model is presented with a set of figure-caption pairs, and its goal is to generate a caption for a given image based on the provided demonstrations. Given a sequential image-captions pairs $S=\{(I_{i},C_{i})\}_{i=1}^{n}$ consisting of $n$ pairs of image $I_{i}$ and the corresponding $C_{i}$, the contextualized image captioning task can be formalized as follows:
$\hat{C_{n}}=\mathcal{M}(I_{1},C_{1},\ldots,I_{n-1},C_{n-1},I_{n},P_{c}).$ (3)
@@
Title Generation This task requires a nuanced understanding of figures and captions to distill essential observations into a high-level summary of the presented results for LVLMs. Specifically, instead of producing the captions for the figures, this task requires the model to connect different figures and corresponding captions to infer the paper title. Let $S=\{(I_{i},C_{i})\}_{i=1}^{m}$ be a sequence of $m$ figure-caption pairs in the extracted paper. Note that $I_{i}$ could be a single figure or a multiple-figure, and we reuse $I_{i}$ for simplicity here. The title generation asks $\mathcal{M}$ to generate the title for the paper given a task prompt $P_{t}$:
$\hat{T}=\mathcal{M}(I_{1},C_{1},\ldots,I_{m},C_{m},P_{t}).$ (4)
@@ 4.2.2 Experimental Settings
 w/ ArXivCap 8.9 15.8 83.3
Bard 3.4 12.2 82.2
GPT-4V 5.7 14.6 83.2
Table 4: Evaluation results of single figure captioning. Grey results are obtained from a 200-sample subset. Despite most LVLMs struggling to produce high-quality captions of scientific figures, training with ArXivCap significantly boosts the performance.
Evaluated Models We select various LVLMs covering different architectures. (1) LVLMs designed for dealing with a single image, BLIP2-OPT-6.7B (Li et al., 2023b) , InstructBLIP-Vicuna7B (Dai et al., 2023), LLaVA-1.5-7B/13B (Liu et al., 2023a). Due to the ability limitation, we only benchmark these models on the single image captioning task; (2) LVLMs capable of handling interleaved text-image inputs, such as OpenFlamingo-9B (Alayrac et al., 2022; Awadalla et al., 2023), IDEFICS-Instruct-9B (Laurençon et al., 2023), Qwen-VL-Chat-7B (Bai et al., 2023b). These models are evaluated on all the tasks we proposed; (3) Proprietary models such as Bard and GPT-4V. Due to the API limit of these models, we evaluate these models on a randomly sampled subset consisting of 200 instances for evaluation, with corresponding scores colored in grey. Details of evaluated models and the task prompts used are provided in Appendix B.
Training Settings To investigate whether in-domain training can enhance the model’s ability, we train the Qwen-VL-Chat-7B for three epochs and adopt the AdamW optimizer (Kingma and Ba, 2015) with a learning rate of 1e-5 and weight decay set to 0.05, following the original setup of Qwen-VL-Chat. We set the maximum figure number in a sample to $4$ to fit the input length limit. The training costs 70 hours with 8 NVIDIA A100 GPUs.
Metrics We adopt BLEU-2 (Papineni et al., 2002), ROUGE-L (Lin, 2004) and BERT-Score (Zhang et al., 2020) as the automatic evaluation metrics. A manual error analysis is performed to supplement the automatic evaluation (§5).
@@ 4.2.3 Results
Results of Single-Figure Captioning The evaluation results for the single figure captioning task are presented in Table 4. Despite achieving almost perfect performance on conventional image captioning tasks such as MSCOCO (Lin et al., 2014), LVLMs encounter challenges when applied to academic figures. Among LVLMs, Qwen-VL-Chat emerges as the top-performing open-source candidate, surpassing its counterparts. Notably, GPT-4V excels in generating captions compared to the competing model, Bard. Moreover, continuous training on our dataset yields a significant performance boost for this task. For instance, fine-tuning results in a notable increase in the BLEU-2 score from 4.4 to 8.9, indicating a promising avenue for enhancing academic figure comprehension through domain-specific training.
Results of Multiple-Figure Captioning As shown in the first block of table 5, similar to single-figure captioning, multiple-image captioning poses a challenge for current open-source LVLMs. For instance, Qwen-VL-Chat achieves only a 3.0 BLEU-2 and a 7.2 ROUGE-L score on this task, considerably lower than its performance in single-figure captioning. In contrast, GPT-4V consistently demonstrates proficiency in both tasks, suggesting a balanced ability to capture semantics across multiple images. Notably, training on our ArXivCap dataset yields more pronounced improvements for this task, culminating in Qwen-VL-Chat even surpassing the performance of the GPT-4V model. This enhancement underscores the pivotal role of our dataset in facilitating LVLMs to enhance reasoning capabilities over multiple images, leading to more effective summarization of scientific figures.
Results of Contextualized Captioning In the middle block of Table 5, we find that IDEFICS-Instruct-9B achieves the best performance on this task. This achievement is largely attributed to its remarkable proficiency in leveraging contextual cues, stemming from its extensive pre-training involving interleaved image-text pairs (Laurençon et al., 2023). Interestingly, fine-tuning on ArXivCap results in marginal performance declines across all metrics, with GPT-4V achieving the lowest scores as well. This phenomenon can be attributed to the tendency of sequential captions to exhibit similar patterns, thereby favoring models that effectively leverage contextual cues. We further perform a more challenging evaluation by randomly shuffling the order of figure-caption pairs in the context. As shown in Table 6, the fine-tuned model demonstrates more robust captioning results under this setting, evidenced by the slight 8% drop on ROUGE-L compared to the 31% of the original model.
@@
Results of Title Generation The results are presented in the last block of Table 5. Notably, the title generation task poses a formidable challenge, evident in the significantly lower overall BLEU-2 score compared to the captioning tasks. This suggests the inherent difficulty in generating precise predictions for paper titles. A contrasting picture emerges when considering the ROUGE-L and BERT-Score metrics, which either closely align or surpass the performance on captioning tasks. This underscores the model’s proficiency in producing semantic-related results given the presented figures. Consistent with the previous two tasks, fine-tuning the model on our dataset yields substantial enhancements for the title generation task. The BLEU-2 score jumps impressively from 2.6 to 6.7, while the ROUGE-L score sees a commendable increase from 15.8 to 23.5. These findings highlight the challenge of title generation for current LVLMs and the effectiveness of our dataset in improving the model’s capability to generate accurate titles.
4.2.4 Analysis
@@
Manual Evaluation We conduct a manual inspection for single-figure captioning results. To ensure a more informed evaluation, we focus on a paper from the CS domain, leveraging our domain knowledge to assess caption quality better. The quality of generated captions is assessed by scrutinizing the figure, the ground-truth caption, the paper title, and the abstract. We categorize captions into the following quality types according to our preliminary inspection: (1) Acceptable, where captions accurately encapsulate the scientific figure’s essence, aligning with the intended information of the ground-truth; (2) Over Simplification, instances where the model oversimplifies content, offering a broad overview while neglecting specific details and nuances present in the ground truth; (3) Recognition Error, where the model inaccurately recognizes and describes key visual and textual elements in the scientific figure, such as colors, numerical values, or textual context; and (4) Contextual Misinterpretation, where the model misinterprets the specific context of the scientific figure, resulting in captions relevant in a generic sense but inaccurate for the given figure. Visualized generated captions of different types are shown in Figure 14 of Appendix C.1. The results of 100 manually examined captions are depicted in Figure 5, revealing that only 16% of captions are deemed acceptable when compared to human-written ones. Among unsatisfactory captions, contextual misinterpretation emerges as the dominant issue, suggesting a need for incorporating more contextual information. Oversimplification is another concern, with generic captions identified. Additionally, 23% and 19% of examined captions suffer from the oversimplification issue and recognition errors in reported numbers/texts in the caption, respectively. The former is attributed to the highly frequent simple caption in the training dataset and the latter issue could be addressed through potential integration with OCR results. Our manual evaluation suggests future efforts may benefit from incorporating additional context clues, such as paper metadata, improving the model’s fundamental perception abilities, and utilizing external information.
@@
Case Study We conducting case studies to illuminate the tuning effects facilitated by our ArXivQA dataset. In the left segment of Figure 6, ArXivQA helps the model accurately answer a question related to the presented bar plot. The right part in Figure 6 demonstrates that ArXivQA can enhance algebraic reasoning abilities. Here, a question involving the derivative of a function is correctly answered, accompanied by a lucid reasoning rationale. Figure 15 in Appendix C.2 highlights a challenging geometry problem where both models generate hallucinated outputs. These illustrative cases collectively affirm the efficacy of our dataset.
@@ 5 Conclusion
Our study has several limitations worth noting. Firstly, our exploration may not encompass the full spectrum of LVLMs due to the rapid evolution of architectures and training methodologies. Nevertheless, we believe our dataset could still be effective for other LVLMs and the findings are generalizable. Secondly, our Multimodal ArXiv primarily sources from ArXiv papers due to its accessibility and open-sourced license, potentially overlooking the diversity of disciplines and data modalities present in scientific literature. Future research could incorporate a broader range of datasets and domains to enrich the coverage of scientific knowledge.

### v3

@@ 4.1.3 Analysis
We investigate how different subject domains affect mathematical reasoning ability using pairs of questions and answers (QA). We focus on six domains with more than 5K samples each. From each domain, we randomly choose a subset of 5K samples to ensure fairness in comparison. We then fine-tune the Qwen-VL-Chat base model using QA pairs from each domain and observe how it affects the model’s accuracy compared to its original state. Figure 4 demonstrates the relative accuracy changes (i.e., Accuracy after Fine-tuningOriginal $\frac{\text{Accuracy after Fine-tuning}}{\text{Original Accuracy}}-1$) after training the model on QA pairs from each domain. Our findings reveal several key points: (i) QA pairs from the Computer Science (CS) domain are highly effective for improving mathematical reasoning ability, achieving a notable 27.09% relative improvement. We attribute this to the compositional nature of the CS area. (ii) The most beneficial domain varies depending on the specific task. For instance, QA pairs from astrophysics domains enhance geometry problem-solving, while those from Condensed Matter improve performance in math word problems. (iii) Most domains hurt the Figure QA task. This suggests that synthetic Figure QA might not be the best benchmark for assessing realistic reasoning ability. These findings underscore the efficacy of generated QA pairs and offer valuable insights for future research, such as adjusting task-specific weights in the dataset accordingly.
@@ 4.2.1 Evaluated Tasks
Four vision-to-text tasks to benchmark LVLMs’ ability to comprehend scientific figures.
Single-Figure Captioning
Similar to the traditional image captioning setup (Lin et al., 2014), single-figure captioning requires the model to generate a caption for the given single figure. The captions generated by the model are expected to encapsulate the nuanced details within these figures, including numbers and mathematical formulas, presenting a unique challenge for models to identify and articulate these elements accurately. Formally, given an image-caption pair $(I,C)$, the LVLM $\mathcal{M}$ is asked to generate the caption given an instruction prompt $P_{s}$ to hint the goal of scientific captioning:
$\hat{C}=\mathcal{M}(I,P_{s}),$
@@
Multiple-Figure Captioning
We introduce a more intricate challenge involving applying reasoning across multiple images. This task, termed Multiple-Figure Captioning, necessitates the model to craft a comprehensive summary caption for subfigures. As exemplified in Figure 2, the model is tasked with generating an overarching caption for two or more subfigures, leveraging visual clues to draw comparisons and formulate summary captions. Formally, given a list of figures $L=\left(I_{1},\ldots,I_{n}\right)$, the model is asked to generate the ground-truth main caption $C$ by considering all the semantics in the figures with a task prompt $P_{m}$:
$\hat{C}=\mathcal{M}(L,P_{m})=\mathcal{M}(I_{1},\ldots,I_{n},P_{m}).$
Contextualized Captioning
Inspired by the evolving in-context learning capabilities of LLMs (Brown et al., 2020b; Dong et al., 2022), we introduce a contextualized captioning task to examine the in-context learning ability of LVLMs. In this task, the model is presented with a set of figure-caption pairs, and its goal is to generate a caption for a given image based on the provided demonstrations. Given a sequential image-captions pairs $S=\{(I_{i},C_{i})\}_{i=1}^{n}$ consisting of $n$ pairs of image $I_{i}$ and the corresponding $C_{i}$, the contextualized image captioning task can be formalized as follows:
$\hat{C_{n}}=\mathcal{M}(I_{1},C_{1},\ldots,I_{n-1},C_{n-1},I_{n},P_{c}).$
@@
Title Generation
This task requires a nuanced understanding of figures and captions to distill essential observations into a high-level summary of the presented results for LVLMs. Specifically, instead of producing the captions for the figures, this task requires the model to connect different figures and corresponding captions to infer the paper title. Let $S=\{(I_{i},C_{i})\}_{i=1}^{m}$ be a sequence of $m$ figure-caption pairs in the extracted paper. Note that $I_{i}$ could be a single figure or a multiple-figure, and we reuse $I_{i}$ for simplicity here. The title generation asks $\mathcal{M}$ to generate the title for the paper given a task prompt $P_{t}$:
$\hat{T}=\mathcal{M}(I_{1},C_{1},\ldots,I_{m},C_{m},P_{t}).$
@@ 4.2.2 Experimental Settings
Dataset
We divide ArXivCap into training and test sets with a 9:1 ratio for evaluation. The test set includes: 161.3K samples for single-figure captioning, 12.8K samples for multiple-figure captioning, 57.2K samples for contextualized captioning, and 57.2K samples for title generation.
Evaluated Models
We select various LVLMs covering different architectures. (1) LVLMs designed for dealing with a single image, BLIP2-OPT-6.7B (Li et al., 2023b) , InstructBLIP-Vicuna7B (Dai et al., 2023), LLaVA-1.5-7B/13B (Liu et al., 2023a). Due to the ability limitation, we only benchmark these models on the single image captioning task; (2) LVLMs capable of handling interleaved text-image inputs, such as OpenFlamingo-9B (Alayrac et al., 2022; Awadalla et al., 2023), IDEFICS-Instruct-9B (Laurençon et al., 2023), Qwen-VL-Chat-7B (Bai et al., 2023b). These models are evaluated on all the tasks we proposed; (3) Proprietary models such as Gemini 1.0 Pro Vision and GPT-4V. Due to the large scale of our test set, we randomly sample a subset consisting of 500 instances for evaluating these two models to reduce costs, with corresponding scores colored in grey. Details of evaluated models and the task prompts used are provided in Appendix B.
Training Settings
To investigate whether in-domain training can enhance the model’s capabilities, we train the Qwen-VL-Chat-7B on ArXivCap using the same setting as in §4.1.1. To fit the input length limit, we set the maximum number of figures per sample to four. The training process takes 70 hours with 8 NVIDIA A100s.
Metrics
BLEU-2 (Papineni et al., 2002), ROUGE-L (Lin, 2004) and BERT-Score (Zhang et al., 2020) are adopted as the automatic evaluation metrics. We also explore using GPT-4 to assist in caption evaluation. Our findings in Appendix B.3 indicate that ROUGE-L and BLEU-2 scores are highly correlated with GPT-4’s annotations. We primarily use these three metrics due to their convenience. A manual error analysis is conducted to supplement the automatic metrics (§4.3).
@@
[…]
Gemini 1.0 Pro Vision 5.6 14.5 82.2
GPT-4V 5.5 14.2 83.3
Table 5: Evaluation results of single figure captioning. Grey results are obtained from a 500-sample subset. Despite most LVLMs struggle to produce high-quality captions of scientific figures, training with ArXivCap significantly boosts the performance.
Model BLEU-2 ROUGE-L BERT-S
Qwen-VL-Chat-7B 4.4 11.1 81.8
   + Title 5.7 13.1 81.6
   + Title and Abstract 6.0 12.7 81.4
[…]
   + Title 12.9 18.6 83.8
   + Title and Abstract 12.7 18.5 83.8
Table 6: Evaluation results of single figure captioning with paper meta information.
@@ 4.2.3 Results
Results of Single-Figure Captioning
The evaluation results for the single-figure captioning task are presented in Table 5. Despite achieving near-perfect performance on conventional image captioning tasks like MSCOCO (Lin et al., 2014), open-source LVLMs, such as LLaVA models, face challenges when applied to academic figures. For closed models, GPT-4V performs comparably with Gemini 1.0 Pro Vision. Furthermore, continuous training on our dataset yields a significant performance boost for this task. For instance, fine-tuning results in a notable increase in the BLEU-2 score from 4.4 to 8.9, indicating a promising avenue for enhancing academic figure comprehension through domain-specific training. We also investigate whether providing additional context information, such as the paper title and abstract, could help models generate better figure captions. As shown in Table 6, adding the title is beneficial evidenced by the boosted scores, while providing abstracts brings negligible gains.
Results of Multiple-Figure Captioning
As shown in the first block of Table 7, similar to single-figure captioning, multiple-image captioning poses a challenge for current open-source LVLMs. For instance, Qwen-VL-Chat achieves only a 3.0 BLEU-2 and a 7.2 ROUGE-L score on this task, considerably lower than its performance in single-figure captioning. In contrast, GPT-4V consistently demonstrates proficiency in both tasks, suggesting a balanced ability to capture semantics across multiple images. Notably, training on our ArXivCap dataset yields more pronounced improvements for this task, culminating in Qwen-VL-Chat even surpassing the performance of the GPT-4V model. This enhancement underscores the pivotal role of our dataset in facilitating LVLMs to enhance reasoning capabilities over multiple images, leading to more effective summarization of scientific figures.
Results of Contextualized Captioning
In the middle block of Table 7, we find that IDEFICS-Instruct-9B achieves the best performance on this task. This achievement is largely attributed to its remarkable proficiency in leveraging contextual cues, stemming from its extensive pre-training involving interleaved image-text pairs (Laurençon et al., 2023). Interestingly, fine-tuning on ArXivCap results in marginal performance declines across all metrics, with GPT-4V achieving the lowest scores as well. This phenomenon can be attributed to the tendency of sequential captions to exhibit similar patterns, thereby favoring models that effectively leverage contextual cues. We perform two more challenging evaluations by (i) providing context pairs from another paper and (ii) randomly shuffling the order of figure-caption pairs in the context. As shown in Table 8, the performance with random contexts degrades significantly, validating our previous hypothesis. Instead, the fine-tuned model demonstrates more robust captioning results under these settings, evidenced by the slight 8% drop on ROUGE-L compared to the 31% of the original model with shuffled context orders.
@@
Results of Title Generation
The results are presented in the last block of Table 7. Notably, the title generation task poses a formidable challenge, evident in the significantly lower overall BLEU-2 score compared to the captioning tasks. This suggests the inherent difficulty in generating precise predictions for paper titles. A contrasting picture emerges when considering the ROUGE-L and BERT-Score metrics, which either closely align or surpass the performance on captioning tasks. This underscores the model’s proficiency in producing semantic-related results given the presented figures. Consistent with the previous two tasks, fine-tuning the model on our dataset yields substantial enhancements for the title generation task. The BLEU-2 score jumps impressively from 2.6 to 6.7, while the ROUGE-L score sees a commendable increase from 15.8 to 23.5. These findings highlight the challenge of title generation for current LVLMs and the effectiveness of our dataset in improving the model’s capability to generate accurate titles.
4.3 Analysis
@@
Manual Evaluation of Generated Captions
We conduct a manual inspection for single-figure captioning results. To ensure a more informed evaluation, we focus on a paper from the CS domain, leveraging our domain knowledge to assess caption quality better. The quality of generated captions is assessed by scrutinizing the figure, the ground-truth caption, the paper title, and the abstract. We categorize captions into the following quality types according to our preliminary inspection: (1) Acceptable, where captions accurately encapsulate the scientific figure’s essence, aligning with the intended information of the ground-truth; (2) Over Simplification, instances where the model oversimplifies content, offering a broad overview while neglecting specific details and nuances present in the ground truth; (3) Recognition Error, where the model inaccurately recognizes and describes key visual and textual elements in the scientific figure, such as colors, numerical values, or textual context; and (4) Contextual Misinterpretation, where the model misinterprets the specific context of the scientific figure, resulting in captions relevant in a generic sense but inaccurate for the given figure. Visualized generated captions of different types are shown in Figure 14 of Appendix C.1. The results of 100 manually examined captions are depicted in Figure 5, revealing that only 16% of captions are deemed acceptable when compared to human-written ones. Among unsatisfactory captions, contextual misinterpretation emerges as the dominant issue, suggesting a need for incorporating more contextual information as suggested in Table 6. Oversimplification is another concern, with generic captions identified. Additionally, 23% and 19% of examined captions suffer from the oversimplification issue and recognition errors in reported numbers/texts in the caption, respectively. The former is attributed to the highly frequent simple caption in the training dataset and the latter issue could be addressed through potential integration with OCR results. Our manual evaluation suggests future efforts may benefit from incorporating additional context clues, such as paper metadata, improving the model’s fundamental perception abilities, and utilizing external information.
Case Study of MathVista
We conduct case studies to illuminate the tuning effects facilitated by our ArXivQA dataset. In the left segment of Figure 6, ArXivQA helps the model accurately answer a question related to the presented bar plot. The right part in Figure 6 demonstrates that ArXivQA can enhance algebraic reasoning abilities. Here, a question involving the derivative of a function is correctly answered, accompanied by a lucid reasoning rationale. Figure 15 in Appendix C.2 highlights a challenging geometry problem where both models generate hallucinated outputs. These illustrative cases collectively affirm the efficacy of our dataset.
@@ 5 Conclusion
Our study has several limitations worth noting. Firstly, our exploration may not encompass the full spectrum of LVLMs due to the rapid evolution of architectures and training methodologies such as parameter-efficient tuning (Hu et al., 2022; Ma et al., 2024). Nevertheless, we believe our dataset could still be effective for other LVLMs and the findings are generalizable. We show that our ArXivQA dataset could also boost LLaVA-series models across scientific understanding benchmarks in Appendix D. Secondly, our Multimodal ArXiv dataset sources from ArXiv papers due to their accessibility and open-source licenses. This approach may overlook the diversity of disciplines and data modalities present in the broader scientific literature. Future research could incorporate a broader range of datasets and domains to enrich the coverage of scientific knowledge, and explore dynamic data selection methods to improve performance and sample efficiency (Li et al., 2021; Chen et al., 2024).
Acknowledgements
We would like to thank all the anonymous reviewers for their insightful comments and suggestions, which helped us improve the quality and clarity of this work. We are particularly grateful to Shuhuai Ren for his valuable feedback in preparing the manuscript. This research was supported in part by the joint research scheme of the National Natural Science Foundation of China (NSFC) and the Research Grants Council (RGC) under grant number N HKU714/21.